### v2

# Fluctuations of Young diagrams for symplectic groups and semiclassical orthogonal polynomials

Abstract.
Consider an $n\times k$ matrix of i.i.d. Bernoulli random numbers with $p=1/2$. Dual RSK algorithm gives a bijection of this matrix to a pair of Young tableaux of conjugate shape, which is manifestation of skew Howe $GL_{n}\times GL_{k}$-duality. Thus the probability measure on zero-ones matrix leads to the probability measure on Young diagrams proportional to the ratio of the dimension of $GL_{n}\times GL_{k}$-representation and the dimension of the exterior algebra $\bigwedge\left(\mathbb{C}^{n}\otimes\mathbb{C}^{k}\right)$.
Similarly, by applying Proctor’s algorithm based on Berele’s modification of the Schensted insertion, we get skew Howe duality for the pairs of groups $Sp_{2n}\times Sp_{2k}$. In the limit when $n,k\to\infty$ $GL$-case is relatively easily studied by use of free-fermionic representation for the correlation kernel. But for the symplectic groups there is no convenient free-fermionic representation. We use Christoffel transformation to obtain the semiclassical orthogonal polynomials for $Sp_{2n}\times Sp_{2k}$ from Krawtchouk polynomials that describe $GL_{2n}\times GL_{2k}$ case. We derive an integral representation for semiclassical polynomials. The study of the asymptotic of this integral representation gives us the description of the limit shapes and fluctuations of the random Young diagrams for symplectic groups.
Introduction
Consider an $n\times k$ matrix of i.i.d. Bernoulli random numbers with $p=1/2$. Dual Robinson-Schensted-Knuth algorithm gives a bijection of this matrix to a pair of semistandard Young tableaux of conjugate shape [1], which is manifestation of skew Howe $GL_{n}\times GL_{k}$-duality [2]. Thus the probability measure on zero-ones matrices leads to the probability measure on Young diagrams proportional to the ratio of the dimension of $GL_{n}\times GL_{k}$-representation and the dimension of the exterior algebra $\bigwedge\left(\mathbb{C}^{n}\otimes\mathbb{C}^{k}\right)$. If $n$ and $k$ go to infinity with the same rate these random diagrams in French notation, rotated by $\frac{\pi}{4}$ and scaled by a factor $\frac{1}{n}$, converge uniformly in probability to a smooth limit shape supported on an interval. This limit shape and fluctuations around it were considered in [3, 4, 5].
Random diagram can be seen as a random point configuration on a lattice, where a point is put in the left boundary of any unit interval, where the upper boundary of diagram decays. That is if diagram $\lambda$ has row lengths $(\lambda_{1},\dots,\lambda_{n})$ the coordinates of the points would be $\{\lambda_{i}+n-i\}_{i=1}^{n}$. Using Weyl dimension formula for irreducible representations of general linear groups it is possible to present this random point ensemble as orthogonal polynomial ensemble with Krawtchouk polynomials. The asymptotic behavior of Krawtchouk polynomials was studied extensively through the years, see [6, 7, 8, 9] and references therein. The corresponding determinatal ensemble was considered in [10, 11, 12, 13].
Skew Howe duality for other pairs of classical Lie groups can be obtained by certain modifications of the insertion algorithm [14]. For example, by applying Proctor’s algorithm [14] based on Berele’s modification [15] of the Schensted insertion [16], we get skew Howe duality for the pairs of groups $Sp_{2n}\times Sp_{2k}$. The algorithm that proves $SO_{2n+1}\times\operatorname{Pin}_{2k}$ duality was proposed by Benkart and Stroomer [17], and for $O_{2n}\times SO_{2k}$ by Okada [18] and Terada [19]. All these algorithms give the bijections between zero-one matrices enumerating basis elements in an exterior algebra and pairs of corresponding modifications of semistandard Young tableaux for classical Lie groups, such as King tableaux for symplectic groups [20] and or Sundaram tableaux for orthogonal groups [21]. The duality leads to the multiplicity-free decomposition of the exterior algebra $\bigwedge\left(\mathbb{C}^{2n}\otimes\mathbb{C}^{k}\right)$ into the direct sum of irreducible representations of dual groups pair. The decomposition gives rise to the probability measure on (symplectic or orthogonal) Young diagrams given by the ratio of dimension of the irreducible component to the dimension of the whole exterior algebra.
Similarly to $GL_{n}\times GL_{k}$-case if we rotate and scale the diagram and take the limit $n,k\to\infty$ with the same rate, we get the uniform convergence to the limit shape. Moreover, the limiting curve is “half” of $GL_{2n}\times GL_{2k}$ curve and is described by the same explicit formula, as demonstrated in [22] by using the dimension formulas and solving the variational problems. The study of the fluctuations in these cases remained an open question.
In the present paper we consider the orthogonal polynomial ensemble that corresponds to the $Sp_{2n}\times Sp_{2k}$-skew Howe duality. As demonstrated in [23], the polynomials are obtained from Krawtchouk polynomials by the “lifting” procedure from [24] which is a variant of the QR-algorithm for the Christoffel transformation of orthognal polynomials [25][26]. The polynomials are therefore discrete semiclassical orthogonal polynomials [27] and their asymptotic behavior was not considered previously. We study various asymptotic regimes for the orthogonal polynomials and obtain the description of local fluctuations for random symplectic Young diagrams.
For brevity we call these polynomials symplectic and they are new in this context. Our main results describe both the exact representation of the correlation kernel in terms of these polynomials and its asymptotic convergence in the bulk regime.
In what follows we introduce the notation:
  • Let $\{\widetilde{K}_{m}(x)\}_{m\geq 0}$ denote the system of monic Krawtchouk orthogonal polynomials on a suitable discrete set, orthogonal with respect to a weight function $W(x)$ (see (2.4)).
  • The modified weight is defined as
    $\widetilde{W}(x)=x^{2}\,W(x),$
    and the corresponding orthonormal system of symplectic polynomials (after applying the Christoffel transformation) is denoted by $\{g_{m}(x)\}_{m\geq 0}$.
  • The correlation kernel of the determinantal point process is given by the Christoffel–Darboux formula
    $\mathcal{K}(u,v)=2\frac{\varkappa_{2n-2}}{\varkappa_{2n}}\frac{g_{2n}(u)g_{2n-2}(v)-g_{2n-2}(u)g_{2n}(v)}{u^{2}-v^{2}}\sqrt{\widetilde{W}(u^{2})\widetilde{W}(v^{2})},$
    where $\varkappa_{m}$ is the leading coefficient of $g_{m}$, and the weight $\widetilde{W}(y)$ is expressed in terms of $y=x^{2}$.
We now state our main theorems.
Theorem 1 (Symplectic polynomials three term relation).
Define the lifted monic polynomials $\{G_{m}(x)\}$ by applying the Christoffel transformation (see (2.7)) corresponding to the change of weight $W(x)$ into $\widetilde{W}(x)=x^{2}\,W(x)$. The orthonormal symplectic polynomials are expressed as $g_{m}(x)=\frac{G_{m}(x)}{\sqrt{\Lambda_{m}}}$ with $\Lambda_{m}$ satisfying nonlinear recurrence relations (2.26),(2.27), (2.28). The coefficients in the three-term recurrence relation satisfied by the symplectic polynomials,
$x\,g_{k}(x)=\tilde{a}_{k}\,g_{k+1}(x)+\tilde{b}_{k}\,g_{k}(x)+\tilde{a}_{k-1}\,g_{k-1}(x),$
are obtained by applying the QR-algorithm to the recurrence coefficients of the original Krawtchouk polynomials (2.19). For $p=1/2$ the recurrence for $\Lambda_{m}$ simplifies to (2.46) and the recurrence for $\{\tilde{a}_{k},\tilde{b}_{k}\}$ is exactly solvable as (2.30) and (2.31). The kernel $\mathcal{K}(x,y)$ represents the orthogonal projection onto the span of $\{g_{0}(x),g_{2}(x),\dots,g_{2n-2}(x)\}$.
Theorem 2 (Sine Kernel Limit).
Let $\mu_{n,k}$ be the measure corresponding to the $Sp_{2n}\times Sp_{2k}$ skew Howe duality, and let $\mathcal{K}(u,v)$ be the associated correlation kernel defined above. Define
$K=2n+2k,\quad\text{and assume }n=HK\text{ with fixed }0<H<\frac{1}{2}.$
Then, in the bulk regime, that is deifned as
$u=\frac{1}{H}\Bigl{(}x+\frac{i}{K}\Bigr{)},\quad v=\frac{1}{H}\Bigl{(}x+\frac{j}{K}\Bigr{)},\quad i,j\in\mathbb{Z},\quad x\in(-1/2,1/2),\quad K\to\infty,$
the kernel converges to the discrete sine kernel in the following sense:
$\lim\limits_{n\to\infty}\frac{\mathcal{K}\left(\frac{1}{H}\big{(}x+\frac{i}{K}\big{)},\frac{1}{H}\big{(}x+\frac{j}{K}\big{)}\right)}{\mathcal{K}(x/H,x/H)}=\frac{\sin(\pi\rho(x)(i-j))}{\pi\rho(x)(i-j)},$
and the limit density is explicitly given by
$\rho(x)=\frac{1}{\pi}\cos^{-1}\!\Biggl{(}\frac{1-4H}{\sqrt{1-4x^{2}}}\Biggr{)}.$
Theorem 1 is established by identifying the Christoffel transformation (2.7) as the mechanism to incorporate the weight modification (i.e., multiplication by $x^{2}$) and by analyzing the induced QR dynamics on the recurrence coefficients. Theorem 2 follows from a steepest descent analysis of the relevant integral representations for the polynomials and shows that the local fluctuations converge to universal sine kernel statistics; the density $\rho(x)$ is consistent with the limit shape analysis in [22].
The paper is organized as follows. In Section 1 we introduce the probability measure on Young diagrams, discuss sampling of random diagrams and recall the definition of the correlation kernel. In Section 2 we write the correlation kernel in terms of the orthogonal polynomials that are obtained from the Krawtchouk polynomials by Christoffel transformation. Then in Section 3 we use the integral representation for the polynomials to study their asymptotic behavior. These results are used to prove the convergence of the correlation kernel to the discrete sine kernel in Section 4. We discuss possible further results in Conclusion.
Acknowledgments
We thank Travis Scrimshaw for reading the first draft and pointing out the deficiencies in it. We thank Anton Dzhamay for useful discussions. The study was carried out with the financial support of the Ministry of Science and Higher Education of the Russian Federation in the framework of a scientific project under agreement No. 075-15-2025-013.
1. Random diagrams for symplectic groups
Consider the symplectic group $Sp_{2n}$ and its first fundamental representation on $\mathbb{C}^{2n}$. The exterior algebra $\bigwedge\left(\mathbb{C}^{2n}\right)$ is a space of a reducible representation of $Sp_{2n}$. The exterior algebra $\bigwedge\left(\mathbb{C}^{2n}\otimes\mathbb{C}^{k}\right)$ can be seen as $k$-th tensor power of this representation $\bigwedge\left(\mathbb{C}^{2n}\otimes\mathbb{C}^{k}\right)=\left(\bigwedge\left(\mathbb{C}^{2n}\right)\right)^{\otimes k}$ and therefore has a natural action of $Sp_{2n}$. On the other hand the same can be done for $Sp_{2k}$, moreover, the actions of $Sp_{2n}$ and $Sp_{2k}$ commute on $\bigwedge\left(\mathbb{C}^{2n}\otimes\mathbb{C}^{k}\right)$ and we have a strongly multiplicity-free decomposition into irreducible representations of $Sp_{2n}\times Sp_{2k}$ [28],[29]
$\bigwedge\left(\mathbb{C}^{2n}\otimes\mathbb{C}^{k}\right)=\bigoplus_{\lambda\subset k^{n}}V_{Sp_{2n}}(\lambda)\otimes V_{Sp_{2k}}(\overline{\lambda}^{\prime}).$ (1.1)
Here $\lambda$ is a Young diagram that is confined inside of $n\times k$ rectangle, $\overline{\lambda}^{\prime}$ is its transposed complement in the rectangle and by $V_{Sp_{2n}}(\lambda)$ we denote the irreducible representation corresponding to the diagram $\lambda$. This is known as skew $(Sp_{2n},Sp_{2k})$ Howe duality [2], [30]. Combinatorially this duality can be understood in terms of King tableaux that generalize semistandard Young tableaux to the symplectic group case.
Let us briefly recall that King tableau [20] for the symplectic group $Sp_{2n}$ is Young diagram $\lambda$ filled in a semistandard way with signed numbers that have order $1<\bar{1}<2<\bar{2}<\dots<n<\bar{n}$. In English convention it means that numbers are strictly increasing along the columns and non-decreasing along the rows. An additional condition is imposed – all the numbers that are $\leq\bar{i}$ should not appear below row number $i$. King tableaux of shape $\lambda$ enumerate the basis vectors in the irreducible representation $V_{Sp_{2n}}(\lambda)$.
Consider the space $\mathbb{C}^{n}$ with the basis vectors $\{e_{i}\}_{i=1}^{n}$. The space $\bigwedge\left(C^{n}\right)$ is $2^{n}$-dimensional and its basis vectors $\{e_{i_{1}}\wedge e_{i_{2}}\wedge\dots\}$ are in one-to-one correspondence to the columns of zeros and ones that encode which vectors appear in the exterior product. Therefore the basis vector of the space $\bigwedge\left(\mathbb{C}^{2n}\otimes\mathbb{C}^{k}\right)=\left(\bigwedge\mathbb{C}^{n}\right)^{\otimes 2k}$ corresponds to the $n\times 2k$ zero-one matrix of $2k$ columns.
The following algorithm based upon Berele insertion [15] was proposed by Proctor in [14] to establish the correspondence between zero-one matrices that encode the basis vectors in the left hand side of the decomposition (1.1) and pairs of King tableaux $(P,Q)$ of complementary shapes $(\lambda,\overline{\lambda}^{\prime})$ inside of $n\times k$ rectangle, that encode the basis vectors in the right hand side of the decomposition. We pair the columns of the zero-one matrix and on each step of the algorithm we consider the pair of columns. On step $r$ the shapes of $P,Q$ form the $n\times r$ rectangle. From two consecutive columns number $12r,2r+$ 1 we produce a sequence of numbers to insert by reading the columns bottom-up. If in row number $i$ we have $(0,0)$ we produce the number $i$, for $(0,1)$ we insert nothing, for $(1,0)$ we insert $\bar{i},i$ and for $(1,1)$ we insert $\bar{i}$. Then we insert numbers one-by-one by using Schensted insertion [16]. If the additional condition is broken by an insertion of number $i$ that is bumping number $\bar{i}$ below the row $i$, we need to erase the box where $\bar{i}$ was. Then we shift this empty box to the boundary as in jeu de taquin (to preserve the semi-standard condition). After the insertion of all the numbers at step $r$, the erased boxes are filled with numbers $\bar{r}$ and adjoined to $Q$. All the empty boxes between the extended tableaux $P$ and $Q$ in $n\times(r+1)$-rectangle are filled with $r$ and adjoined to $Q$ as well. Then we proceed to the step number $(r+1)$.
Besides the combinatorial proof of the decomposition (1.1), Proctor’s algorithm gives us an effective way to sample random Young diagrams by generating random zero-one matrices with some measure and then pushing this measure to Young diagrams.
By considering the dimension of the components in the decomposition (1.1), we get the identity
$2^{2nk}=\sum_{\lambda\subset k^{n}}\dim V_{Sp_{2n}}(\lambda)\cdot\dim V_{Sp_{2k}}(\overline{\lambda}^{\prime}),$
which can be used to introduce the probability measure on the diagrams that appear in the decomposition. The probability of a diagram is the ratio of the dimension of the irreducible $Sp_{2n}\times Sp_{2k}$-representation corresponding to the diagram to the dimension of the whole exterior algebra, this measure corresponds to the uniform measure on zero-one matrices
$\mu_{n,k}(\lambda)=2^{-2nk}\dim V_{Sp_{2n}}(\lambda)\cdot\dim V_{Sp_{2k}}(\overline{\lambda}^{\prime}).$ (1.2)
It is possible to write this measure explicitly, as was demonstrated in [22]. To do so introduce the coordinates $a_{i}=\lambda_{i}+n-i+1$. These coordinates can be interpreted as the coordinates of unit intervals where the upper boundary of Young diagram $\lambda$, rotated by $\frac{\pi}{4}$ (“Russian convention”), is decreasing. The diagram then can be identified with a particle configuration $\left\{a_{i}\right\}$ on the integer lattice. For $\dim V_{Sp_{2n}}(\lambda)$ we can use Weyl dimension formula. The problem is to write the dimension of $V_{Sp_{2k}}(\overline{\lambda}^{\prime})$ in terms of the coordinates $\{a_{i}\}$, which are introduced for the diagram $\lambda$. It can be done by representing King tableau as a set of non-intersecting paths, then switching to the dual paths and using Lindström–Gessel–Viennot lemma to express the number of non-intersecting paths configurations as a determinant [31]. Then Dodgson condensation leads to the explicit formula
$\mu_{n,k}(\left\{a_{i}\right\})=\frac{2^{2n(1-k)}}{\prod_{i<j}(j-i)(2n+2-i-j)}\prod_{i<j}(a_{i}^{2}-a_{j}^{2})^{2}\prod_{\ell=1}^{n}a_{\ell}^{2}\frac{(2k-1+2\ell)!}{(k+n+a_{\ell})!(k+n-a_{\ell})!}.$ (1.3)
By this derivation the measure is determinantal, therefore the measure and all of the correlation functions can be expressed as the determinants of the correlation kernel
$\mu_{n,k}(\left\{a_{i}\right\})=\det\left[\mathcal{K}(a_{i},a_{j})\right]_{i,j=1}^{n}.$ (1.4)
In particular the density of particles is just a one-point correlation function
$\rho(a)=\mathcal{K}(a,a).$ (1.5)
We are interested in the limit behavior of the probability measure when $n$ and $k$ go to infinity with the same rate. We assume that the size of the box of the diagram goes to zero with the same rate by introducing the coordinate $u=\frac{a}{n}$.
In [22] we have demonstrated that random Young diagrams, rotated by $\frac{\pi}{4}$ and considered as 1-Lipschitz functions, converge uniformly in probability to the limit curve $\Omega(u)$ that has an explicit formula for the limit density $\mathbb{P}(\sup_{u}|\lambda(u)-\Omega(u)|>\varepsilon)\to 0$ as $n\to\infty$. In [23] the same was demonstrated for a very similar problem in $(SO_{2n+1},\operatorname{Pin}_{2k})$ skew Howe duality. To do so we have written the measure in the exponential form, used Stirling approximation for the factorials and solved the variational problem. But we have not considered the local fluctuations around the limit shape. Nevertheless in [23] in a similar problem we have demonstrated that the probability measure can be written in terms of a semiclassical family of orthogonal polynomials. This representation was then used to prove Central Limit Theorem for the global fluctuations. Below we use the polynomials to write the correlation kernel in the Christoffel-Darboux form and demonstrate its convergence to the discrete sine kernel in the bulk by analyzing the asymptotic behavior of the orthogonal polynomials.
In our case the measure (1.3) contains a Vandermonde determinant written in terms of squares of variables $a_{i}$. Introduce the variables $y_{i}=\frac{a_{i}^{2}}{n^{2}}$, then measure takes the determinantal form
$\mu_{n,k}(y_{1},\dots,y_{n})=\frac{1}{Z_{n,k}}\prod_{i<j}(y_{i}-y_{j})^{2}\cdot\prod_{\ell=1}^{n}y_{\ell}\binom{2n+2k}{n+k-n\sqrt{y_{\ell}}},$ (1.6)
where we have rewritten the weight in the binomial form and collected the terms that do not depend on $y_{i}$ into the normalization constant:
$Z_{n,k}^{-1}=\frac{2^{2n(1-k)}n^{2n^{2}}}{\prod\limits_{i<j}^{n}(j-i)(2n+2-i-j)}\prod\limits_{l=1}^{n}\frac{(2k-1+2l)!}{(2k+2n)!}$
Denote the weight by $\widetilde{W}(y)$
$\widetilde{W}(y)=y\;\binom{2n+2k}{n+k-n\sqrt{y}},$ (1.7)
and introduce the discrete orthonormal polynomials $p_{m}(y)$ on a quadratic lattice
$\left\{\frac{i^{2}}{n^{2}}|i=0,1,\dots,n+k\right\}$
$\sum_{i=0}^{n+k}p_{m}\left(\frac{i^{2}}{n^{2}}\right)p_{\ell}\left(\frac{i^{2}}{n^{2}}\right)\widetilde{W}\left(\frac{i^{2}}{n^{2}}\right)=\delta_{m,\ell}.$ (1.8)
According to (1.7), the weight $\widetilde{W}(u^{2})$ is the binomial coefficient multiplied by $u^{2}$. The polynomials, that are orthogonal with the binomial weight are the Krawtchouk polynomials. The polynomials $p_{m}$ are related to the Krawtchouk polynomials using the “lifting” procedure from [24] which is a variant of the QR-algorithm for the Christoffel transformation of orthognal polynomials [25][26] as demonstrated below.
Introduce Christoffel-Darboux correlation kernel as
$\mathcal{K}(y_{i},y_{j})=\sqrt{\widetilde{W}(y_{i})\widetilde{W}(y_{j})}\sum_{\ell=0}^{n-1}p_{\ell}(y_{i})\;p_{\ell}(y_{j}),$ (1.9)
the measure is then rewritten as its determinant
$\mu_{n,k}(y_{1},\dots,y_{n})=\frac{1}{Z_{n,k}}\det\left[\mathcal{K}(y_{i},y_{j})\right]_{i,j=1}^{n}=\frac{1}{Z_{n,k}}\det\left(\sqrt{\widetilde{W}(y_{i})\widetilde{W}(y_{j})}\sum_{l=0}^{n-1}p_{l}(y_{i})\;p_{l}(y_{j})\right)_{i,j=1}^{n}.$ (1.10)
We are interested in the asymptotic behavior of the probability measure as $n,k$ go to infinity with the same rate. Therefore we need to study the asymptotics of the polynomials $\{p_{m}\}$. First note, that the polynomials $\{p_{m}(y)\}$ with $y=u^{2}$ are just even polynomials from the set of polynomials $\{g_{m}(u)\}$ that are orthonormal with respect to the weight $\widetilde{W}(u^{2})$
$\sum_{i=-(n+k)}^{n+k}g_{m}\left(\frac{i}{n}\right)g_{\ell}\left(\frac{i}{n}\right)\widetilde{W}\left(\frac{i^{2}}{n^{2}}\right)=\delta_{m,\ell}.$ (1.11)
Since weight $\widetilde{W}(u^{2})$ is even as a function of $u$ and since each value of $y$ appears in (1.11) twice, we have
$p_{m}\left(u^{2}\right)=\sqrt{2}g_{2m}(u),$ (1.12)
for all m we can write the kernel in terms of the polynomials $\{g_{2m}\}$ using Christoffel-Darboux formula as
$\mathcal{K}(u,v)=2\frac{\varkappa_{2n-2}}{\varkappa_{2n}}\frac{g_{2n}(u)g_{2n-2}(v)-g_{2n-2}(u)g_{2n}(v)}{u^{2}-v^{2}}\sqrt{\widetilde{W}(u^{2})\widetilde{W}(v^{2})},$ (1.13)
where $\varkappa_{m}$ is the coefficient of the $m$-th power of $u$ in $g_{m}(u)$, with variables $u,v$ living on a linear lattice $\left\{-\frac{n+k}{n},-\frac{n+k-1}{n},\dots,\frac{n+k-1}{n},\frac{n+k}{n}\right\}$.
We study the asymptotics $n,k\to\infty$ such that their ratio tends to constant. Introduce $K=2n+2k$, and $H$ such that $n=HK$, then we are interested in the limit $K\to\infty$ with $H$ being constant. Bulk asymptotic regime corresponds to $u=\frac{1}{H}\left(x+\frac{i^{\prime}}{n}\right)$, $v=\frac{1}{H}\left(x+\frac{j^{\prime}}{n}\right)$ with $i^{\prime},j^{\prime}\in\mathbb{Z}$. In the next Section we express the polynomials in correlation kernel in terms of Krawtchouk polynomials and then study their asymptotic behavior.
2. A family of semiclassical orthogonal polynomials
2.1. Symplectic polynomials
Non-normalised Krawtchouk polynomials $\{K_{m}\}$ satisfy the orthogonality relation
$\sum\limits_{a=0}^{K}\binom{K}{a}p^{a}(1-p)^{K-a}K_{m}(a,p,K)K_{n}(a,p,K)=\frac{(-1)^{n}n!}{(-K)_{l}}\left(\frac{1-p}{p}\right)^{n}\delta_{mn}=\binom{K}{n}^{-1}\left(\frac{1-p}{p}\right)^{n}\delta_{nm},$
where $p\in(0,1)$ and we use the following definition of the Pochhammer symbol
$(-K)_{m}=\prod\limits_{i=0}^{m}(-K+i-1)=(-1)^{m}m!\binom{K}{m}.$
There is a well known representation of the Krawtchouk polynomials as hypergeometric functions [32]
$K_{m}(a,p,K)=\,_{2}F_{1}\left(-m,-a,-K;\frac{1}{p}\right).$ (2.1)
The monic version of the $K_{m}(a,p,K)$ is defined as (9.11.4 in [32])
$P_{m}(a,p,K)=(-K)_{m}p^{m}K_{m}(a,p,K).$ (2.2)
So the monic Krawtchouk polynomial $P_{m}(a,p,K)$ now has the following hypergeometric representation (from the (2.1))
$P_{m}(a,p,K)=(-1)^{m}m!p^{m}\binom{K}{m}K_{m}(a,p,K)=(-1)^{m}m!p^{m}\binom{K}{m}\,_{2}F_{1}\left(-m,-a,-K;\frac{1}{p}\right)$ (2.3)
To arrive to the polynomials $g_{m}(u)$ we first shift the $a$ variable to $a\to K/2+nu$, where the $u$ is the new variable on the symmetric lattice. Denote the monic polynomial with rescaled variable by $\tilde{K}_{m}(u)\equiv\tilde{K}_{m}(u,p,K)$. We immediately find their hypergeometric representation
$\widetilde{K}_{m}(u)$ $=\frac{1}{n^{m}}P_{m}(K/2+nu,p,K)=(-1)^{m}\frac{p^{m}}{n^{m}}m!\binom{K}{m}K_{m}(K/2+nu,p,K)=$
$=(-1)^{m}\frac{p^{m}}{n^{m}}m!\binom{K}{m}\,_{2}F_{1}\left(-m,-K/2-nu,-K;\frac{1}{p}\right).$ (2.4)
The orthogonality relation for polynomials $\widetilde{K}_{m}$ is
$\sum\limits_{i=-K/2}^{K/2}\widetilde{K}_{l}(i/n)\widetilde{K}_{m}(i/n)W(i/n)=n^{-2m}\kappa^{-2}_{m}\delta_{lm},$ (2.5)
where $\kappa_{m}=\frac{(-1)^{m}}{m!}\binom{K}{m}^{-1/2}(p(1-p))^{-m/2}$, and we have the relation between the weights $W$ for Krawtchouk polynomials for the variable on the symmetric lattice $u\in\{-\frac{K}{2n},-\frac{K}{2n}+1/n,...,\frac{K}{2n}\}$ and weight $\widetilde{W}$ for $\{g_{m}(u)\}$,
$u^{2}W(u)=u^{2}\binom{K}{K/2+nu}p^{K/2-nu}(1-p)^{K/2+nu}=\widetilde{W}(u^{2}).$ (2.6)
Since the weight $\widetilde{W}(u^{2})$, defined by (1.7) differs from the weight of Krawthouk polynomials $W$ with $K=2n+2k$ only by a factor $u^{2}$, monic orthogonal polynomials $G_{m}$ are expressed in terms of the polynomials (2.4) by the formula (Christoffel transformation, see [33], sect. 2.7, Theorem 2.7.1)
$G_{m}(u)=\frac{1}{C_{m,2}u^{2}}\left|\begin{matrix}\widetilde{K}_{m}(0)&\widetilde{K}_{m+1}(0)&\widetilde{K}_{m+2}(0)\\ \left(\widetilde{K}_{m}\right)^{\prime}(0)&\left(\widetilde{K}_{m+1}\right)^{\prime}(0)&\left(\widetilde{K}_{m+2}\right)^{\prime}(0)\\ \widetilde{K}_{m}(u)&\widetilde{K}_{m+1}(u)&\widetilde{K}_{m+2}(u)\end{matrix}\right|,$ (2.7)
where
$C_{m,2}=\left|\begin{matrix}\widetilde{K}_{m}(0)&\widetilde{K}_{m+1}(0)\\ \left(\widetilde{K}_{m}\right)^{\prime}(0)&\left(\widetilde{K}_{m+1}\right)^{\prime}(0)\end{matrix}\right|.$ (2.8)
We are interested only in the even polynomials $g_{2m}$. As monic polynomials $\widetilde{K}_{m}$ are defined on a symmetric interval, we have $\widetilde{K}_{2l+1}(0)=0$ and $\widetilde{K}_{2l}^{\prime}(0)=0$. Then $C_{2m,2}=\widetilde{K}_{2m}(0)\widetilde{K}_{2m+1}^{\prime}(0)$ and
$G_{2m}(u)$ $=\frac{1}{u^{2}}\left(\widetilde{K}_{2m+2}(u)-\frac{\widetilde{K}_{2m+2}(0)}{\widetilde{K}_{2m}(0)}\widetilde{K}_{2m}(u)\right).$ (2.9)
By the same considerations, $C_{2m+1,2}=-\widetilde{K}_{2m+2}(0)\widetilde{K}_{2m+1}^{\prime}(0)$ and
$G_{2m+1}(u)$ $=\frac{1}{u^{2}}\left(\widetilde{K}_{2m+3}(u)-\frac{\widetilde{K}_{2m+3}^{\prime}(0)}{\widetilde{K}_{2m+1}^{\prime}(0)}\widetilde{K}_{2m+1}(u)\right).$ (2.10)
We denote the orthogonality relation as
$\sum\limits_{i=-K/2}^{K/2}(i/n)^{2}W(i/n)G_{m}(i/n)G_{k}(i/n)=\Lambda_{m}\delta_{mk},$ (2.11)
where $\Lambda_{m}=||G_{m}||$ can be computed by the recurrence (see next section). The monic symplectic polynomials $\{G_{m}(u)\}$ have the following hypergeometric realization for even $m$
$G_{m}(u)=\frac{(-1)^{m}p^{m+2}(m+2)!}{u^{2}n^{m+2}}\binom{K}{m+2}\bigg{(}\,_{2}F_{1}\left(-m-2,-K/2-nu,-K;\frac{1}{p}\right)-$
$-\frac{\,{}_{2}F_{1}\left(-m-2,-K/2,-K;\frac{1}{p}\right)}{\,{}_{2}F_{1}\left(-m,-K/2,-K;\frac{1}{p}\right)}\,_{2}F_{1}\left(-m,-K/2-nu,-K;\frac{1}{p}\right)\bigg{)},$ (2.12)
It is also useful to note that this hypergeometric representation significantly simplifies for $p=1/2$ because of the Gauss’s formula $\frac{\,{}_{2}F_{1}\left(-m-2,-K/2,-K;2\right)}{\,{}_{2}F_{1}\left(-m,-K/2,-K;2\right)}=\frac{m+1}{(-K+m+1)}$ (see Appendix),
$G_{m}(u)=\frac{1}{u^{2}\kappa_{m+2}n^{m+2}}\binom{K}{m+2}^{1/2}\bigg{(}\,_{2}F_{1}\left(-m-2,-K/2-nu,-K;2\right)+$
$+\frac{(m+1)}{(K-m-1)}\,_{2}F_{1}\left(-m,-K/2-nu,-K;2\right)\bigg{)}.$ (2.13)
Remark 1.
We define the symplectic polynomials $\{g_{m}(u)\}$ as a Christoffel transformation of the polynomials $\{\widetilde{K}_{m}(u)\}$ (see (2.7)) for any $p\in(0,1)$ from (2.4). All methods used in this work are defined for $p\in(0,1)$ unless otherwise specified, however, all basic calculations are performed only in the case of $p=1/2$. The reason for this is that symplectic polynomials for $p\neq 1/2$ do not have a clear probabilistic interpretation as (1.10).
2.2. QR algorithm and three term relation’s coefficients
Using the results from [24], one can derive a three-term recurrence relation for symplectic polynomials. This can be achieved by employing the fact that the Christoffel transformation provides a method for constructing orthogonal polynomials with respect to the measure $x^{2}W(x)dx$ from those that are orthogonal with respect to the positive Borel measure $W(x)dx$.
Introducing $\pi_{m}(x)=n^{m}\kappa_{m}\widetilde{K}_{m}(x)$, one can check that the polynomials $\{\pi_{m}\}$ satisfy the conditions
[…]
with the measure
$d\mu(x)=dxW(x)\sum\limits_{i=-K/2}^{K/2}\delta(x-i/n),$
that is automatically normalized,
[…]
Their three term recurrence relation is given by Jacobi matrix
$x\pi_{m}(x)=a_{m+1}\pi_{m+1}(x)+b_{m}\pi_{m}(x)+a_{m}\pi_{m-1}(x),$ (2.15)
with the starting conditions $\pi_{-1}(x)=0,\,\pi_{0}(x)=1.$ It is obvious that the Christoffel transformation from the $\pi_{m}(x)$ is the same as multiplying the result of Christoffel transformation from $\tilde{K}_{m}(x)$ by $n^{m+2}\kappa_{m+2}$, so we recall
$x^{2}n^{m+2}\kappa_{m+2}G_{m}(x)=\pi_{m+2}(x)-\frac{\pi_{m+2}(0)}{\pi_{m}(0)}\pi_{m}(x).$ (2.16)
We also need to compute the second moment
[…]
$=\frac{1}{n^{2}}\sum\limits_{j=0}^{K}(j-K/2)^{2}\binom{K}{j}p^{j}(1-p)^{K-j}=\frac{1}{4n^{2}}K\left(K(1-2p)^{2}-4(p-1)p\right).$ (2.17)
Now we can determine the coefficients of the three term relation for the polynomials $\{\pi_{m}(x)\}$ from three term relation coefficients $\{\tilde{\alpha}_{m},\tilde{\beta}_{m}\}$ of the monic polynomials $\{\widetilde{K}_{m}\}$ (see [32])
$x\widetilde{K}_{m}(x)=\widetilde{K}_{m+1}(x)+\tilde{\alpha}_{m}\widetilde{K}_{m}(x)+\tilde{\beta}_{m}\widetilde{K}_{m-1}(x)=$ (2.18)
$=\widetilde{K}_{m+1}(x)+[p(K-m)+m(1-p)-K/2]\widetilde{K}_{m}(x)+\frac{m(K-m+1)}{n^{2}}p(1-p)\widetilde{K}_{m-1}(x).$
So we get the Jacobi matrix coefficients for (2.15)
$a_{m}=\frac{\kappa_{m-1}}{n\kappa_{m}}=-\frac{\big{[}m(K-m+1)(p(1-p))\big{]}^{1/2}}{n}=-\sqrt{\tilde{\beta}_{m}},$
$b_{m}=\left[p(K-m)+m(1-p)-K/2\right]=\tilde{\alpha}_{m}.$
For $p=1/2$ the weight function $W(x)$ becomes symmetric and $b_{m}=0$ as expected.
Following the Theorem 3.3 (see [24]), consider the Jacobi matrix of the three term relation coefficient for the orthonormal polynomials with the measure $\mu$. Denote by $\omega$ the natural mapping from the set of all normalized, real, positive Borel measures $\mathcal{B}$ to the set of all singly-infinite, symmetric, tridiagonal irreducible matrices $\mathcal{J}$
$\omega:\mathcal{B}\to\mathcal{J},$
that takes every member of $\mathcal{B}$ to the corresponding Jacobi matrix. By the Favard theorem [34] this mapping is invertible if and only if the underlying Hamburger problem is determinate [35]. By the inverse mapping we mean $\omega^{-1}:\mathcal{J}\to\mathcal{B}^{*}$, where $\mathcal{B}^{*}$ is the quotient by the equivalence classes of the measures with identical moment sequences. In our case the measure $d\mu(x)$ defined above has the discrete and finite support, which means that the moment problem for $\text{Im}(\omega)$ is always determinate and the equivalence class of the element $\omega^{-1}(J)$ consists of a single element (see [36], [37], [38]). Then the QR algorithm gives the reversible arrows in the diagram
${J}$${\widehat{J}}$${d\mu}$${d\tilde{\mu}}$$\scriptstyle{\tau}$$\scriptstyle{\omega}$$\scriptstyle{\rho}$$\scriptstyle{\omega}$
where the map $\tau$ is a step of $QR$ algorithm, as applied to a square matrix $J$, replaces it by the product $\widehat{J}=RQ$, where $J=QR$, Q is orthogonal and R is upper triangular.
We can identify the Cristoffel transformation of $\pi_{m}(x)$ with the normalized measure transformation, that is defined by the commutative diagram above
$\omega^{-1}\circ\tau\circ\omega=\rho:d\mu(x)\to d\tilde{\mu}(x)=\frac{x^{2}d\mu(x)}{\Omega_{2}(K,p,n)}.$
$QR$ factorization of Jacobi matrix $J$ is its decomposition into the product $J=Q_{1}^{T}Q_{2}^{T}...Q_{n-1}^{T}R$, where $Q_{i}$ are block-diagonal antisymmetric matrices, which turn the subdiagonal elements to zero (Givens rotation). For example, letting $r_{0}=\sqrt{a_{1}^{2}+b_{0}^{2}}$, we take $Q_{1}$ to be identity except that the upper left $2\times 2$ block is
$Q_{1}\big{|}_{(e_{1},e_{2})}=\begin{bmatrix}b_{0}/r_{0}&a_{1}/r_{0}\\ -a_{1}/r_{0}&b_{1}/r_{0}\end{bmatrix}.$
Then the result of the first Givens rotation acting on the initial matrix $J$ is
$Q_{1}J\;=\;\begin{bmatrix}r_{0}&(a_{1}b_{0}+a_{1}b_{1})/r_{0}&(a_{1}a_{2})/r_{0}&\\ &(b_{0}b_{1}-a_{1}^{2})/r_{0}&(b_{0}a_{2})/r_{0}&0\\ &a_{2}&b_{2}&a_{3}&\ddots\\ &&a_{3}&b_{3}&\ddots\\ &&&\ddots&\ddots\end{bmatrix}=\begin{bmatrix}r_{0}&s_{1}&(a_{1}a_{2})/r_{0}&\\ &b_{1}^{*}&a_{2}^{*}&0\\ &a_{2}&b_{2}&a_{3}&\ddots\\ &&a_{3}&b_{3}&\ddots\\ &&&\ddots&\ddots\end{bmatrix}.$
The trailing coefficients $b_{i}^{*}$ and $a_{i+1}^{*}$, obtained by applying $i$ successive Givens rotations to $J$, serve as the parameters for the subsequent rotation $Q_{i+1}$. The nontrivial part of $Q_{i+1}$, in its $(k+1)$st and $(k+2)$nd rows and columns, has the form
$Q_{i+1}\big{|}_{(e_{i+1},e_{i+2})}=\begin{bmatrix}b_{i}^{*}/r_{i}&a_{i+1}/r_{i}\\ -a_{i+1}/r_{i}&b_{i}^{*}/r_{i}\end{bmatrix}.$
The resulting upper triangular matrix $R$ is defined by diagonal elements $r_{0},r_{1},\dots$ and superdiagonal elements $s_{1},s_{2},\dots$, respectively, where
$a^{*}_{1}=a_{1},\quad b_{0}^{*}=b_{0},$
$a_{k}^{*}=\frac{a_{k}b^{*}_{k-2}}{r_{k-2}},\quad b^{*}_{k}=\frac{b^{*}_{k-1}b_{k}-a^{*}_{k}a_{k}}{r_{k-1}},$
$r^{2}_{k}=a^{2}_{k+1}+b^{*2}_{k},\quad s_{k}=\frac{a_{k}^{*}b^{*}_{k-1}+a_{k}b_{k}}{r_{k-1}}=\frac{a_{k}b_{k-2}^{*}b_{k-1}^{*}}{r_{k-2}r_{k-1}}+\frac{a_{k}b_{k}}{r_{k-1}}.$
The elements above the superdiagonal in the upper-triangular matrix $R$ are not involved in the QR-transformed matrix $\widehat{J}$ and are irrelevant for our purposes. The elements of the QR-transformed Jacobi matrix $\widehat{J}=RQ$ are defined by
$\hat{a}_{k}=\frac{a_{k}r_{k}}{r_{k-1}},\quad\hat{b}_{k}=\frac{b_{k-1}^{*}b_{k}^{*}}{r_{k-1}}+\frac{a_{k+1}s_{k+1}}{r_{k}},$ (2.19)
with $\hat{b}_{0}=b_{0}+\frac{a_{1}s_{1}}{r_{0}}$.
The three-term relation for modified polynomials is then
$x\hat{g}_{m}(x)=\hat{a}_{m+1}\hat{g}_{m+1}(x)+\hat{b}_{m}\hat{g}_{m}(x)+\hat{a}_{m}\hat{g}_{m-1}(x),$ (2.20)
where $\hat{g}_{m}(x)$ satisfy orthogonal relation
$\sum\limits_{i=-K/2}^{K/2}\hat{g}_{m}(i/n)\hat{g}_{l}(i/n)(i/n)^{2}\frac{W(i/n)}{\Omega_{2}(K,p,n)}=\delta_{ml}.$ (2.21)
There is the direct computation of the recurense relations of the normalization $\Lambda_{m}$, which appears in (2.11) as the norm of the symplectic polynomials. Let us define the three term relation
$xG_{m}(x)=G_{m+1}(x)+\alpha_{m}G_{m}(x)+\beta_{m}G_{m-1}(x),$ (2.22)
then
$\sum_{x\in\mathfrak{X}}G_{m}(x)G_{l}(x)x^{2}W(x)=\delta_{ml}\Lambda_{m},\quad\mathfrak{X}=\bigg{\{}-\frac{K}{2n},-\frac{K}{2n}+\frac{1}{n},...,\frac{K}{2n}\bigg{\}}.$ (2.23)
Denoting $n^{-2m}\kappa_{m}^{-2}=L_{m}$ in (2.5), $S_{2l}=-\widetilde{K}_{2l+2}(0)/\widetilde{K}_{2l}(0)$ in (2.9), $S_{2l+1}=-\widetilde{K}_{2l+3}^{\prime}(0)/\widetilde{K}_{2l+1}^{\prime}(0)$ in (2.10), one can compute
$\sum_{x\in\mathfrak{X}}G_{m}(x)G_{l}(x)x^{4}W(x)=\sum_{x\in\mathfrak{X}}(\widetilde{K}_{m+2}(x)+S_{m}\widetilde{K}_{m}(x))(\widetilde{K}_{l+2}(x)+S_{l}\widetilde{K}_{l}(x))W(x)=$
$=\delta_{m,l}L_{m+2}+\delta_{m,l}S_{m}^{2}L_{m}+\delta_{m,l+2}S_{m}L_{m}+\delta_{m+2,l}S_{l}L_{l}.$ (2.24)
On the other hand, using (2.22), we get
$\sum_{x\in\mathfrak{X}}G_{m}(x)G_{l}(x)x^{4}W(x)=$
$=\sum_{x\in\mathfrak{X}}(G_{m+1}(x)+\alpha_{m}G_{m}(x)+\beta_{m}G_{m-1}(x))(G_{l+1}(x)+\alpha_{l}G_{l}(x)+\beta_{l}G_{l-1}(x))x^{2}W(x)=$
$=\delta_{m,l}\Lambda_{m+1}+\delta_{m+2,l}\beta_{l}\Lambda_{m+1}+\delta_{m,l+2}\beta_{m}\Lambda_{m-1}+\delta_{m,l}\beta_{m}^{2}\Lambda_{m-1}+\delta_{ml}\alpha_{m}^{2}\Lambda_{m}+\delta_{m+1,l}\alpha_{l}\Lambda_{m+1}+$
$+\delta_{m-1,l}\beta_{m}\alpha_{l}\Lambda_{m-1}+\delta_{m,l+1}\alpha_{m}\Lambda_{m}+\delta_{m,l-1}\alpha_{m}\beta_{l}\Lambda_{m}.$ (2.25)
Comparing these two results we get the recurrences
$(m=l+2)$ $:\quad S_{m}L_{m}=\beta_{m}\Lambda_{m-1},$ (2.26)
$(m=l)$ $:\quad L_{m+2}+S_{m}^{2}L_{m}=\Lambda_{m+1}+\beta_{m}^{2}\Lambda_{m-1}+\alpha_{m}^{2}\Lambda_{m},$ (2.27)
$(m=l+1)$ $:\quad 0=\beta_{m}\alpha_{m-1}\Lambda_{m-1}+\alpha_{m}\Lambda_{m}.$ (2.28)
Now, comparing (2.21) with (2.11), we get
$g_{m}(x)=\frac{\hat{g}_{m}(x)}{\sqrt{\Omega_{2}}},$
then
$\tilde{a}_{m}^{2}=\hat{a}_{m}^{2}=\beta_{m},\quad\tilde{b}_{m}=\hat{b}_{m}=\alpha_{m}.$ (2.29)
It is not so hard to see that in the case $p=1/2$ all $b_{m}=0$ and the Givens rotatation’s parameters become
$a^{*}_{1}=a_{1}=\frac{(-1)}{2HK^{1/2}}=\frac{x}{\pi_{1}(x)},\quad b_{0}^{*}=b_{0}=0=b_{m},$
$a_{k}^{*}=\frac{a_{k}b^{*}_{k-2}}{r_{k-2}}=\frac{-a_{k}a_{k-2}^{*}a_{k-2}}{r_{k-2}r_{k-3}},\quad b^{*}_{k}=\frac{-a^{*}_{k}a_{k}}{r_{k-1}}.$
And now we can express diagonal and superdiagonal elements of the matrix $R$ as
$r^{2}_{k}=a^{2}_{k+1}+\frac{(a^{*}_{k}a_{k})^{2}}{r_{k-1}^{2}},\quad s_{k}=\frac{a_{k}^{*}b^{*}_{k-1}}{r_{k-1}}=\frac{-a_{k}^{*}a_{k-1}^{*}a_{k-1}}{r_{k-1}r_{k-2}},$
$r_{0}^{2}=a_{1}^{2},\quad r_{1}^{2}=a_{2}^{2}+\frac{a_{1}^{4}}{r_{0}^{2}}=a_{2}^{2}+a_{1}^{2},\quad r_{2}^{2}=a_{3}^{2},$
$a_{2}^{*}=0,\quad a_{3}^{*}=-\frac{a_{3}a_{1}^{2}}{r_{0}r_{1}}=-\frac{a_{3}r_{0}}{r_{1}},\quad a_{4}^{*}=0.$
Here $a_{0}$ should be the constant on which nothing depends. Then we can compute
$a_{2m}^{*}=0,\quad a_{2m+1}^{*}=-\frac{a_{2m+1}a_{2m-1}^{*}a_{2m-1}}{r_{2m-1}r_{2m-2}}=-\frac{a_{2m+1}a_{2m-1}^{*}}{r_{2m-1}},$
$b_{2m}^{*}=0,\quad b_{2m+1}^{*}=-a^{*}_{2m+1},\quad s_{k}=0,\;k>0.$
The recursion on the diagonal coefficients of $R$ gives
$r_{2m}^{2}=a_{2m+1}^{2},$ (2.30)
$r_{2m+1}^{2}=a_{2m+2}^{2}+\frac{(a_{2m+1}^{*}a_{2m+1})^{2}}{r_{2m}^{2}}=a_{2m+2}^{2}+a_{2m+1}^{*2}=a_{2m+2}^{2}+\frac{(a_{2m+1}a_{2m-1}a_{2m-3}^{*})^{2}}{r^{2}_{2m-1}r_{2m-3}^{2}}=$
$=a_{2m+2}^{2}+a_{2m+1}^{2}\prod\limits_{j=1}^{m}\frac{a_{2j-1}^{2}}{r_{2j-1}^{2}}=a_{2m+2}^{2}+\frac{\prod\limits_{j=1}^{m+1}a_{2j-1}^{2}}{\sum\limits_{\ell=-1}^{m-1}\left(\prod\limits_{k=0}^{\ell}a_{2m-2k}^{2}\right)\left(\prod\limits_{j=1}^{m-\ell-1}a_{2j-1}^{2}\right)}.$ (2.31)
Here we have expressed the product directly, assuming that $\prod\limits_{k=0}^{-1}(..)=1$,
$\prod\limits_{j=1}^{m}r^{2}_{2j-1}=\left(a_{2m}^{2}+a_{2m-1}^{2}\prod\limits_{j=1}^{m-1}\frac{a_{2j-1}^{2}}{r_{2j-1}^{2}}\right)\prod\limits_{j=1}^{m-1}r^{2}_{2j-1}=\prod\limits_{j=1}^{m}a_{2j-1}^{2}+a_{2m}^{2}\prod\limits_{j=1}^{m-1}r_{2j-1}^{2}=$
$=\prod\limits_{j=1}^{m}a_{2j-1}^{2}+a_{2m}^{2}\prod\limits_{j=1}^{m-1}a_{2j-1}^{2}+...+\left(\prod_{k=0}^{\ell}a_{2m-2k}^{2}\right)\left(\prod_{j=1}^{m-\ell-1}a_{2j-1}^{2}\right)+...+\prod_{k=0}^{m-1}a_{2m-2k}^{2}=$
$=\sum\limits_{\ell=-1}^{m-1}\left(\prod_{k=0}^{\ell}a_{2m-2k}^{2}\right)\left(\prod_{j=1}^{m-\ell-1}a_{2j-1}^{2}\right).$ (2.32)
Finally we express the QR transformed coefficients
$\hat{b}_{0}=0,\quad\hat{b}_{k}=0\;(k>0),$ (2.33)
$\hat{a}^{2}_{2m}=\frac{a^{2}_{2m}a^{2}_{2m+1}}{r^{2}_{2m-1}},\quad\hat{a}^{2}_{2n+1}=r^{2}_{2n+1},$ (2.34)
which ends the proof of Theorem 1.
The expression for $r_{2m+1}^{2}$, more suitable for asymptotic analysis is given by the next lemma.
Lemma 1.
Let $K,m,n\in\mathbb{Z}_{\geq 0}$. Suppose the sequence $\{r_{2m+1}^{2}\}$ satisfies the recurrence
$r_{2m+1}^{2}=a_{2m+2}^{2}+a_{2m+1}^{2}\prod\limits_{j=1}^{m}\frac{a_{2j-1}^{2}}{r_{2j-1}^{2}}.$
Then $r_{2m+1}^{2}$ can be expressed as
$\frac{r_{2m+1}^{2}}{a_{2m+2}^{2}}=\frac{\mathcal{R}^{[2m+1]}_{m}}{\mathcal{S}^{[2m+1]}_{m}},$ (2.35)
where $\mathcal{R}_{k}^{[2m+1]},\,\mathcal{S}_{k}^{[2m+1]}$ satisfy three term reccurenses
$\mathcal{R}_{k}^{[2m+1]}=B_{k}^{[2m+1]}\mathcal{R}^{[2m+1]}_{k-1}+A_{k}^{[2m+1]}\mathcal{R}_{k-2}^{[2m+1]},$ (2.36)
$\mathcal{S}_{k}^{[2m+1]}=B_{k}^{[2m+1]}\mathcal{S}_{k-1}^{[2m+1]}+A_{k}^{[2m+1]}\mathcal{S}_{k-2}^{[2m+1]},$ (2.37)
with starting conditions
$\mathcal{R}_{-1}^{[2m+1]}=1,\quad\mathcal{S}^{[2m+1]}_{-1}=0,$
$\mathcal{R}_{0}^{[2m+1]}=B_{0}^{[2m+1]},\quad\mathcal{S}_{0}^{[2m+1]}=1,$
and the coefficients that are rational functions of $12m+$ 1
$A_{k}^{[2m+1]}=-\frac{a_{2m+1-2(k-1)}^{2}}{a_{2m+2-2(k-1)}^{2}},\quad B_{k}^{[2m+1]}=1+\frac{a_{2m+1-2k}^{2}}{a_{2m+2-2k}^{2}}=1-A_{k+1}.$
Proof.
Using the recursion
$r_{2m+1}^{2}=a_{2m+2}^{2}+a_{2m+1}^{2}\left(\prod\limits_{j=1}^{m-1}\frac{a_{2j-1}^{2}}{r_{2j-1}^{2}}\right)\frac{a_{2m-1}^{2}}{r_{2m-1}^{2}}=a_{2m+2}^{2}+\frac{a_{2m+1}^{2}a_{2m-1}^{2}}{r_{2m-1}^{2}}\frac{(r_{2m-1}^{2}-a_{2m}^{2})}{a_{2m-1}^{2}}=$
$=a_{2m+2}^{2}+a_{2m+1}^{2}\bigg{(}1-\frac{a_{2m}^{2}}{r_{2m-1}^{2}}\bigg{)},$
we get
$\frac{r_{2m+1}^{2}}{a_{2m+2}^{2}}=1+\frac{a_{2m+1}^{2}}{a_{2m+2}^{2}}\left(1-\frac{1}{r_{2m-1}^{2}/a_{2m}^{2}}\right),\quad\frac{r^{2}_{1}}{a_{2}^{2}}=1+\frac{a_{1}^{2}}{a_{2}^{2}}.$
This can be expressed as a convergent of the continued fractions [^1] In this sense, formula (2.32) is analogous to the Euler’s continued fraction formula $\gamma_{0}+\gamma_{0}\gamma_{1}+\gamma_{0}\gamma_{1}\gamma_{2}+\cdots+\gamma_{0}\gamma_{1}\gamma_{2}\cdots\gamma_{n}=\frac{\gamma_{0}}{1+}\,\frac{-\gamma_{1}}{1+\gamma_{1}+}\,\cfrac{-\gamma_{2}}{1+\gamma_{2}+}\cdots\frac{-\gamma_{n}}{1+\gamma_{n}}.$
$\frac{r_{2m+1}^{2}}{a_{2m+2}^{2}}=\bigg{(}1+\frac{a_{2m+1}^{2}}{a_{2m+2}^{2}}\bigg{)}-\frac{a_{2m+1}^{2}/a_{2m+2}^{2}}{r_{2m-1}^{2}/a_{2m}^{2}}=\bigg{(}1+\frac{a_{2m+1}^{2}}{a_{2m+2}^{2}}\bigg{)}-\frac{a_{2m+1}^{2}/a_{2m+2}^{2}}{\left(1+\dfrac{a_{2m-1}^{2}}{a_{2m}^{2}}\right)-\dfrac{a_{2m-1}^{2}/a_{2m}^{2}}{r_{2m-3}^{2}/a_{2m-2}^{2}}}=$
$=B_{0}^{[2m+1]}+\frac{A_{1}^{[2m+1]}}{B_{1}^{[2m+1]}+}\;\frac{A_{2}^{[2m+1]}}{B_{2}^{[2m+1]}+}...\frac{A_{m}^{[2m+1]}}{B_{m}^{[2m+1]}},$
where
$A_{k}^{[2m+1]}=-\frac{a_{2m+1-2(k-1)}^{2}}{a_{2m+2-2(k-1)}^{2}},\quad B_{k}^{[2m+1]}=1+\frac{a_{2m+1-2k}^{2}}{a_{2m+2-2k}^{2}}=1-A_{k+1}^{[2m+1]}.$
Then there exist the unique solutions of three term reccurence for $k\geq 1$ (see, for example, [39])
$\mathcal{R}_{k}^{[2m+1]}=B_{k}^{[2m+1]}\mathcal{R}^{[2m+1]}_{k-1}+A_{k}^{[2m+1]}\mathcal{R}_{k-2}^{[2m+1]},$
$\mathcal{S}_{k}^{[2m+1]}=B_{k}^{[2m+1]}\mathcal{S}_{k-1}^{[2m+1]}+A_{k}^{[2m+1]}\mathcal{S}_{k-2}^{[2m+1]},$
with starting conditions
$\mathcal{R}_{-1}^{[2m+1]}=1,\quad\mathcal{S}^{[2m+1]}_{-1}=0,$
$\mathcal{R}_{0}^{[2m+1]}=B_{0}^{[2m+1]},\quad\mathcal{S}_{0}^{[2m+1]}=1,$
such that
$\frac{r_{2m+1}^{2}}{a_{2m+2}^{2}}=\frac{\mathcal{R}^{[2m+1]}_{m}}{\mathcal{S}^{[2m+1]}_{m}}.$
Corollary 1.
Now we note that in the double-scaling limit $n=HK,m=\frac{\lambda}{2}K,\;,\lambda\in(0,1),\;H\in(0,1/2),\;K\to\infty$ one can find for $k+1\leq m$
$\mathcal{S}_{k}^{[2m+1]}=\mathcal{S}^{[2m+1]}_{k-1}-A_{k+1}^{[2m+1]}\mathcal{S}_{k-1}^{[2m+1]}+A_{k}^{[2m+1]}\mathcal{S}_{k-2}^{[2m+1]}=\mathcal{S}_{k-2}^{[2m+1]}+O(1/K).$ (2.38)
For odd m, we extend the recursion (2.36),(2.37) with the conditions
$A_{k}^{[2m+1]}=\begin{cases}-\frac{a_{2m+1-2(k-1)}^{2}}{a_{2m+2-2(k-1)}^{2}},\quad k\leq m,\\ 0,\quad k>m,\end{cases}$
$B_{k}^{[2m+1]}=\begin{cases}1-A_{k+1},\quad k<m,\\ 1+a_{1}^{2}/a_{2}^{2},\quad k=m,\\ 1,\quad k>m\end{cases}$
so that the three-term relations stabilize at $k\geq m+1$ and we can write
$\frac{r_{2m+1}^{2}}{a_{2m+2}^{2}}=\frac{\mathcal{R}^{[2m+1]}_{m}}{\mathcal{S}^{[2m+1]}_{m}}=\frac{\mathcal{R}^{[2m+1]}_{m+1}}{\mathcal{S}^{[2m+1]}_{m+1}}=\frac{\mathcal{R}^{[2m+1]}_{0}}{\mathcal{S}^{[2m+1]}_{0}}+O(1/K)=1+\frac{a_{2m+1}^{2}}{a_{2m+2}^{2}}+O(1/K).$ (2.39)
Then, the asymptotics of the coefficients of three term relation (2.20) for $p=1/2$ are
$\hat{b}_{0}=0,\quad\hat{b}_{k}=0\;(k>0),$ (2.40)
$\hat{a}^{2}_{2m}=\frac{a^{2}_{2m}a^{2}_{2m+1}}{r^{2}_{2m-1}}=\frac{a_{2m}^{2}a_{2m+1}^{2}}{a_{2m}^{2}+a_{2m-1}^{2}}+O(1/K),$ (2.41)
$\hat{a}^{2}_{2m+1}=r^{2}_{2m+1}=a_{2m+1}^{2}+a_{2m+2}^{2}+O(1/K).$ (2.42)
In order to find the next terms of the asymptotic expansion in the limit under consideration, it is useful to note that the coefficients of the three-term recurrence relations (2.36) and (2.37) are slowly varying functions [^2] A function $f(k)$ is referred to as slowly varying if $f(k)=F(\eta k)$, where $\eta$ is a small parameter and $F$ is a smooth function. of the natural parameter:
$A_{k+1}^{[2m+1]}-A_{k}^{[2m+1]}=-\left(\frac{a_{2m+1-2k}^{2}}{a_{2m+2-2k}^{2}}-\frac{a_{2m+1-2(k-1)}^{2}}{a_{2m+2-2(k-1)}^{2}}\right)=O(1/K^{2}).$ (2.43)
One may then employ the WKB method [40] to construct the asymptotic solution of the three-term recurrence relations (2.36) and (2.37).
Remark 2.
An alternative nonlinear recursion for determining the coefficients $\tilde{a},\tilde{b}$ when $p\in(0,1)$ follows from equations (2.26), (2.27), and (2.28). Eliminating $\Lambda_{m}$ from these equations, we obtain
$L_{m+2}+S_{m}^{2}L_{m}$ $=\frac{S_{m+2}L_{m+2}}{\beta_{m+2}}+\beta_{m}S_{m}L_{m}+\alpha_{m}^{2}\frac{S_{m+1}L_{m+1}}{\beta_{m+1}},$ (2.44)
$S_{m}L_{m}\alpha_{m-1}$ $=-\alpha_{m}\frac{S_{m+1}L_{m+1}}{\beta_{m+1}}.$ (2.45)
The first of these equations is nonlinear and of second order. It remains an open question whether this system can be reduced to two coupled nonlinear first-order recurrences, for example by methods analogous to those in [41] and [42].
In the special case $p=1/2$, one finds $\alpha_{m}=0$, and the remaining recurrence
$L_{m+2}+S_{m}^{2}L_{m}=\frac{S_{m+2}L_{m+2}}{\beta_{m+2}}+\beta_{m}S_{m}L_{m}=\Lambda_{m+1}+\frac{S_{m}^{2}L_{m}^{2}}{\Lambda_{m-1}},$ (2.46)
becomes first order on the subsequences $\beta_{2m}$ and $\beta_{2m+1}$, recovering directly the analogue of the recursion from Lemma 1.
For illustrative purposes we present first several monic Krawtchouk polynomials $\{\widetilde{K}_{m}(x)\}$ and monic symplectic polynomials $\{G_{m}(x)\}$ for $p=1/2$ in Table 1
$\begin{array}[]{|l|l|}\hline\cr\widetilde{K}_{m}(x)&G_{m}(x)\\ \hline\cr 1&1\\ x&x\\ x^{2}-\frac{K}{4\,n^{2}}&x^{2}-\frac{3\,K+2}{4\,n^{2}}\\ x^{3}-\frac{1}{4}\,x{\left(\frac{3\,K-2}{n^{2}}\right)}&x^{3}-\frac{1}{4}x\frac{\left(15K^{2}-30K+16\right)}{(3K-2)n^{2}}\\ x^{4}-\frac{1}{2}\,x^{2}{\left(\frac{3\,K-4}{n^{2}}\right)}+\frac{3\,K^{2}-6\,K}{16\,n^{4}}&x^{4}-\frac{5}{2}x^{2}\frac{(K-2)}{n^{2}}+\frac{15K^{2}-50K+24}{16n^{4}}\\ x^{5}-\frac{5}{2}\,x^{3}{\left(\frac{K-2}{n^{2}}\right)}+\frac{1}{16}\,x{\left(\frac{15\,K^{2}-50\,K+24}{n^{4}}\right)}&x^{5}-\frac{5}{2}x^{3}\frac{21K^{3}-126K^{2}+224K-96}{\left(15K^{2}-50K+24\right)n^{2}}+\\ &\quad+x\frac{525K^{4}-4200K^{3}+11340K^{2}-11840K+4416}{16\left(15K^{2}-50K+24\right)n^{4}}\\ x^{6}-\frac{5}{4}\,x^{4}{\left(\frac{3\,K-8}{n^{2}}\right)}+\frac{1}{16}\,x^{2}{\left(\frac{45\,K^{2}-210\,K+184}{n^{4}}\right)}-&x^{6}+x^{4}\frac{70-21K}{4n^{2}}+\frac{7}{16}x^{2}\frac{15K^{2}-90K+112}{n^{4}}-\\ \quad-\frac{15\,K^{3}+90\,K^{2}-120\,K}{64\,n^{6}}&\quad-\frac{3}{64}\frac{35K^{3}-280K^{2}+588K-240}{n^{6}}\\ \hline\cr\end{array}$
Table 1. Monic Krawtchouk polynomials and monic symplectic polynomials for $p=1/2$.
3. Asymptotics of the polynomials
In investigating the asymptotic properties of the Christoffel–Darboux kernel (1.13), we are led, via the analysis of equation (2.9), to the necessity of determining the asymptotic behavior of the polynomials $\{\widetilde{K}_{m}\}$.
3.1. Double-scaling asymptotic for Krawtchouk polynomials
In this section we analyse the double-scaling limit of the Krawtchouk polynomials. Specifically, we examine the limit in which the density of nodes tends to infinity while the lattice size increases simultaneously. This corresponds to the asymptotic regime given by:
$n=HK,\quad m=gK+s\sqrt{K}+l,\quad K\to\infty,$
where $g,s,l$ are fixed parameters, $10<g<$ 1, and the critical point is located at $g=\frac{1}{2}$.
Numerous studies have explored the asymptotic behavior of the Krawtchouk polynomials $\widetilde{K}_{m}(i/n)$ using methods such as steepest descent and the Riemann-Hilbert approach (see [6], [8], [9]). In our analysis, we aim to extend these computations to the regime where the lattice variable $i$ remains unbounded in the double-scaling limit.
To ensure that $i$ is not constrained, we introduce a new variable $\mu$ defined as
$i=\frac{K}{2}(1-2\mu),\quad i\in[-K/2,K/2],$
where $10<\mu<$ 1 and $\mu\neq\frac{1}{2}$. Now let us consider on the asymptotic expansion of
$\widetilde{K}_{m}(i/n)=(-1)^{m}\frac{m!}{n^{m}}\int_{\Gamma}(1-(1-p)z)^{K/2-i}(1+pz)^{K/2+i}z^{-m}\frac{\mathrm{dz}}{2\pi iz},$ (3.1)
where the contour $\Gamma$ is a positive oriented closed unity loop that encloses 0, $|z|=1$. Changing the variable $t=1/(1-(1-p)z)$ one can find the integral
$I=\int_{\Gamma}(1-(1-p)z)^{K/2-i}(1+pz)^{K/2+i}z^{-m}\frac{\mathrm{dz}}{2\pi iz}=$
$=-(-1)^{m+1}\frac{(1-p)^{m-i-K/2}}{2\pi i}\int_{\Gamma^{\prime}}\frac{t^{s\sqrt{K}+l-1}dt}{(1-t)^{s\sqrt{K}+l+1}}\exp\left\{K\ln{\frac{(t-p)^{1-\mu}}{(1-t)^{g}}}t^{g-1}\right\},$
where the contour $\Gamma^{\prime}$ is a smooth negative -oriented circle around 1. Now we can use saddle-point method to analyze the asymptotic behavior of the integral (see [43])
[…]
$\theta_{\text{SDP}}^{j}=-\frac{\alpha_{j}}{2}\pm\frac{\pi}{2},\quad\alpha_{j}=\arg(f^{\prime\prime}(t_{0}^{j})),$
with
$G(t)=\frac{t^{s\sqrt{K}+l-1}}{(1-t)^{s\sqrt{K}+l+1}},$
$f(t)=(1-\mu)\ln(t-p)+(g-1)\ln t-g\ln(1-t),$
where the $\theta_{\text{SDP}}^{j}=\arg\left(\frac{dt}{d\rho}\right)_{\rho=0}$ is the tangent angle of the steepest descent path in the saddle point $t_{0}^{j}$, $\rho$ is the parameter along the minimax contour and the sign in the expression of $\theta_{\text{SDP}}^{j}$ defined by the steepest descent path orientation in the saddle points.
The equation $f^{\prime}(t)=0$ has two solutions, which we will index by $\pm$,
$t_{0}^{\pm}(p)=\frac{1}{2\mu}\left(\pm\sqrt{(-g+\mu+p)^{2}-4\mu p(1-g)}-g+\mu+p\right).$
Limiting our consideration with the case $p=1/2$, we can simplify this expression to
$t_{0}^{\pm}(1/2)\equiv t_{0}^{\pm}=\frac{\sqrt{2\mu(1-g)}}{2\mu}e^{\pm i\chi},$
where for $(\mu-1/2)^{2}\leq g(1-g)$  [^3] Here and below we use the next definition of Heaviside $\Theta$-function $\Theta(x)=\begin{cases}0,\quad x\leq 0,\\ 1,\quad x>0.\end{cases}$ ,
$\chi=\tan^{-1}\left(\frac{\sqrt{(1-g)g-\left(\mu-1/2\right)^{2}}}{\mu+1/2-g}\right)+\pi\Theta(g-1/2-\mu),$
and in the case $(\mu-1/2)^{2}>g(1-g)$ we have two real roots
$t_{0}^{\pm}=\frac{1}{2\mu}\left(\pm\sqrt{(\mu-1/2)^{2}-g(1-g)}+(\mu+1/2-g)\right),$
In fact it is easy to see that in general
$t_{0}^{\pm}=\frac{1}{2\mu}\bigg{(}\pm\sqrt{(g+\mu-1/2)^{2}-2g\mu}-((g-\mu)-1/2)\bigg{)}.$
So exactly up to factor $\frac{1}{2\mu}$ and changing $\mu\to p$, roots of the action in the considered limit the same as the roots of the generalized Krawtchouk polynomials with $p\neq 1/2$, described above. Starting from the region of two complex-conjugate roots $1/2-\sqrt{\mu(1-\mu)}<g<1/2+\sqrt{\mu(1-\mu)}$ for a fixed $\mu$ (for a fixed $g$, this region is defined by the inequality $1/2-\sqrt{g(1-g)}<\mu<1/2+\sqrt{g(1-g)}$)  [^4] This is also a direct consequence of the self-duality of the Krawtchouk polynomials., let us choose the parametrization $g=1/2+d\sqrt{\mu(1-\mu)},\;d\in[-1,1]$. Then one can easily define two branches of an algebraic rational plane curve $t_{0}(d;\mu)=x(d;\mu)+iy_{\pm}(d;\mu)$
$t_{0}^{\pm}(d;\mu)=\frac{1}{2\mu}\bigg{(}\mu-d\sqrt{\mu(1-\mu)}\pm i\sqrt{\mu(1-\mu)(1-d^{2})}\bigg{)}.$
It is not so hard to see that $t_{0}(d;\mu)$ for fixed $\mu$ form a circle with the equation
$\left(x-\frac{1}{2}\right)^{2}+y_{\pm}^{2}=\frac{(1-\mu)}{4\mu}.$
When we take the limit $\mu\to 0$ we have the circle with infinite radius. For $\mu\to 1$ this circle disappeares symmetrically at $(1/2,0)$ by hyperbollic scaling. There is critical point $(1,0)$ when $\mu=1/2,\,g=0$ and the unique real root goes to center of the integration. The points of the intersection of this circle and line $y=0$ are
$x=\frac{1}{2}\pm\frac{1}{2}\sqrt{\frac{1-\mu}{\mu}}.$
In the case of two real roots we have two real sections instead of a circle
$(g-1/2)^{2}>\mu(1-\mu)\;\Longleftrightarrow\;g(1-g)<(\mu-1/2)^{2},$
so this case is realized for the parametric inequality
$g>1/2+\sqrt{\mu(\mu-1)},\quad g<1/2-\sqrt{\mu(1-\mu)}.$
Using the parametrization
$g(1-g)=\tilde{d}(\mu-1/2)^{2},\quad 0<\tilde{d}<1,$
$g=1/2\pm\sqrt{1/4-\tilde{d}(\mu-1/2)^{2}},$ (3.3)
where the sign is consistent with the sign in the inequalities above, we can immediately write
$t_{0}^{+}(-,\tilde{d})=\frac{1}{2\mu}\bigg{(}(\mu-1/2)\sqrt{1-\tilde{d}}+\mu+\sqrt{1/4-\tilde{d}(\mu-1/2)^{2}}\bigg{)},$
$t_{0}^{-}(-,\tilde{d})=\frac{1}{2\mu}\bigg{(}-(\mu-1/2)\sqrt{1-\tilde{d}}+\mu+\sqrt{1/4-\tilde{d}(\mu-1/2)^{2}}\bigg{)},$
for minus sign, and
$t_{0}^{+}(+,\tilde{d})=\frac{1}{2\mu}\bigg{(}(\mu-1/2)\sqrt{1-\tilde{d}}+\mu-\sqrt{1/4-\tilde{d}(\mu-1/2)^{2}}\bigg{)},$
$t_{0}^{-}(+,\tilde{d})=\frac{1}{2\mu}\bigg{(}-(\mu-1/2)\sqrt{1-\tilde{d}}+\mu-\sqrt{1/4-\tilde{d}(\mu-1/2)^{2}}\bigg{)},$
for the plus sign, where the signs under the brackets denote the choice of the sign in (3.3). These real sections touch each other in pairs at the points $x=\frac{1}{2}\pm\frac{1}{2}\sqrt{\frac{1-\mu}{\mu}}$, when $\tilde{d}=1$.
Let us focus on the case of two complex conjugate roots. One can express all main values from the equation (3.2)
$\operatorname{Re}(f(t_{0}^{\pm}))=\frac{1}{2}\left(\ln\left(\frac{1-\mu}{2(1-g)}\right)+2g\tanh^{-1}(1-2g)+\mu\ln\left(\frac{4\mu}{1-\mu}\right)\right),$ (3.4)
$\operatorname{Im}(f(t_{0}^{\pm}))=\pm\bigg{(}(1-\mu)\gamma+(g-1)\chi-g\tau\bigg{)}=\pm\bigg{(}\zeta_{1}+g\zeta_{2}-\mu\gamma\bigg{)},$ (3.5)
$G(t_{0}^{\pm})=\frac{1}{(t_{0}^{\pm})^{2}}\left(\frac{t_{0}^{\pm}}{1-t_{0}^{\pm}}\right)^{s\sqrt{K}+l+1}=\frac{2\mu}{g}\left(\sqrt{\frac{1-g}{g}}\right)^{s\sqrt{K}+l-1}e^{\pm i\big{[}\zeta_{2}(s\sqrt{K}+l-1)-2\tau\big{]}},$ (3.6)
$\arg f^{\prime\prime}(t_{0}^{\pm})=\pm\bigg{[}\tan^{-1}\left(\frac{(g-1/2)\left((g-1/2)^{2}-\frac{3}{4}\mu+\frac{\mu^{2}}{2}\right)}{\sqrt{g(1-g)-(\mu-1/2)^{2}}\cdot\left(\frac{\mu}{4}-(g-1/2)^{2}\right)}\right)+\sigma+$
$\quad+\pi\cdot\Theta\left(\frac{1}{2}-\frac{\sqrt{\mu}}{2}-g\right)-\pi\cdot\Theta\left(g-\left(\frac{1}{2}+\frac{\sqrt{\mu}}{2}\right)\right)\bigg{]}.$ (3.7)
in the compact form, through the functions $\tau,\gamma,\zeta_{1},\zeta_{2},\sigma$, which are defined as follows. Functions $\tau$ and $\gamma$ are naturally defined as in the case of function $\chi$ as the arguments of the shifted roots of the function $f$
$1-t_{0}^{\pm}=\sqrt{\frac{g}{2\mu}}e^{\pm i\tau},\quad$ $\tau=-\tan^{-1}\left(\frac{\sqrt{(1-g)g-\left(\mu-1/2\right)^{2}}}{g+\mu-1/2}\right)-\pi\Theta(1/2-g-\mu),$
$t_{0}^{\pm}-1/2=\frac{\sqrt{\mu(1-\mu)}}{2\mu}e^{\pm i\gamma},\quad$ $\gamma=\tan^{-1}\left(\frac{\sqrt{(1-g)g-\left(\mu-1/2\right)^{2}}}{1/2-g}\right)+\pi\Theta(g-1/2).$
Functions $\zeta_{1}$ and $\zeta_{2}$ are the angles of these shifts relative to $\chi=\pm\arg(t_{0}^{\pm}(1/2))$
$\zeta_{1}=\gamma-\chi=\tan^{-1}\left(\frac{\sqrt{g(1-g)-(\mu-1/2)^{2}}}{3/2-g-\mu}\right)+\pi\Theta(\mu+g-3/2),$
$\zeta_{2}=\chi-\tau=\tan^{-1}\left(\frac{\sqrt{g(1-g)-(\mu-1/2)^{2}}}{\mu-1/2}\right)+\pi\Theta(1/2-\mu).$
The $\sigma$ is defined from the denominator of $-\pi\leq\arg f^{\prime\prime}(t_{0}^{\pm})\leq\pi$ as
$\sigma=-2\pi\cdot\Theta\left(g-\frac{1}{2}+\frac{\sqrt{(2\mu-3)\mu(1-\mu)}}{2\sqrt{\mu-1}}\right)+2\pi\cdot\Theta\left(g-\frac{1}{2}-\frac{\sqrt{(2\mu-3)\mu(1-\mu)}}{2\sqrt{\mu-1}}\right),$
Now we need to define a Steepest Descent Path (SDP). First of all, lets define the lines of constant phase
$\operatorname{Im}[f(t)-f(t_{0}^{\pm})]=\arg\left(\frac{(t-1/2)^{1-\mu}}{(t_{0}^{\pm}-1/2)^{1-\mu}}\frac{t^{g-1}}{(t_{0}^{\pm})^{g-1}}\frac{(1-t_{0}^{\pm})^{g}}{(1-t)^{g}}\right)=$
$=(1-\mu)\arg(t-1/2)+(g-1)\arg(t)-g\arg(1-t)\mp\bigg{(}\zeta_{1}+g\zeta_{2}-\mu\gamma\bigg{)}=0,$ (3.8)
for $t\in\mathbb{C},\,\mu,g\in\mathbb{R},\,\mu\in(0,1),\,g\in(1/2-\sqrt{\mu(1-\mu)},1/2+\sqrt{\mu(1-\mu})$, and
$\operatorname{Re}[f(t)-f(t_{0}^{\pm})]\leq 0,$
$\operatorname{Re}[f(t_{0}^{\pm})]=\frac{1}{2}\left(\ln\left(\frac{1-\mu}{2(1-g)}\right)+2g\tanh^{-1}(1-2g)+\mu\ln\left(\frac{4\mu}{1-\mu}\right)\right).$
The solutions of these equations are symmetric relative to the reflection $t\to\overline{t}$, it follows from the logarithmicity of the action $f(t)$. By the symmetry of the SDP branches
$\theta^{+}_{\text{SDP}}=-\theta_{\text{SDP}}^{-}+\pi,$
where
$\theta_{\text{SDP}}^{-}=-\frac{\arg(f^{\prime\prime}(t^{-}_{0}))}{2}+\frac{\pi}{2},$ (3.9)
and the sign of the second part is based on the following consideration. Fix $\mu$ to some arbitrary value from its domain and denote the values $L:=1/2-\sqrt{(\mu-1)\mu},\,R:=1/2+\sqrt{(\mu-1)\mu}$. From the SDP’s described above, in the case $L<g<1/2$ we have
$\lim_{g\to 1/2-}\theta_{SDP}^{-}=\pi/2<\theta_{SDP}^{-}<\lim_{g\to L+}\theta_{SDP}^{-}=\pi,$
$\lim_{g\to 1/2-}\alpha_{-}=\frac{1}{2}\lim_{g\to 1/2-}\arg(f^{\prime\prime}(t_{0}^{-}))=\arg(\mu^{2})=0.$
Then in this region of $g$ it follows
$\theta_{\text{SDP}}^{-}=-\frac{\arg(f^{\prime\prime}(t_{0}^{-}))}{2}+\frac{\pi}{2},\quad L<g<1/2.$
In the case $1/2<g<R$ we have
$\lim_{g\to R-}\theta_{SDP}^{-}=0<\theta_{SDP}^{-}<\lim_{g\to 1/2+}\theta_{SDP}^{-}=\pi/2,$
so
$\theta_{\text{SDP}}^{-}=-\frac{\arg(f^{\prime\prime}(t_{0}^{-}))}{2}+\frac{\pi}{2},\quad 1/2<g<R$
$\theta_{\text{SDP}}^{-}=\pi/2,\quad g=1/2.$
We can now construct the asymptotic of the initial integral
[…]
$\sim(-1)^{m}(1/2)^{m-K/2}\frac{(\sqrt{1-g})^{m-1/2}}{(\sqrt{g})^{m+1/2}}\left(\sqrt{\frac{\mu}{1-\mu}}\right)^{K/2-j}\left(\sqrt{\frac{1-\mu}{1-g}}\right)^{K}\sqrt{\frac{2}{\pi K}}\left(\frac{(1-\mu)\mu}{(1-g)g-\left(\mu-\frac{1}{2}\right)^{2}}\right)^{1/4}\cdot$
$\cdot\sin(\theta_{\text{SDP}}^{-}-\hat{\delta}(j;m)),$
where:
$\hat{\delta}(j;m)=(m-1)\zeta_{2}+K\zeta_{1}+(j-K/2)\gamma-2\tau.$
Using the fact that $K$ is even we can continue the simplification of $\hat{\delta}$ by defining new variables
$\hat{\zeta}_{1}=\tan^{-1}\left(\frac{\sqrt{g(1-g)-(\mu-1/2)^{2}}}{3/2-g-\mu}\right),$
$\hat{\tau}=\tan^{-1}\left(\frac{\sqrt{(1-g)g-\left(\mu-1/2\right)^{2}}}{g+\mu-1/2}\right),$
in which the definition of $\hat{\delta}$ depends only on the location of $\mu,g$ relative to $1/2$ value
$\hat{\delta}(j;m)=(m-1)\zeta_{2}+K\hat{\zeta_{1}}-K\mu\gamma+2\hat{\tau},$ (3.10)
and after further simplification of the integral asymptotic we have
[…]
$\sim\frac{(\sqrt{1-g})^{m-K-1/2}(\sqrt{1-\mu})^{K/2+j+1/2}}{(\sqrt{g})^{m+1/2}(\sqrt{\mu})^{j-K/2-1/2}}\sqrt{\frac{2}{\pi K}}\frac{(-1)^{m}(1/2)^{m-K/2}}{\left((1-g)g-\left(\mu-\frac{1}{2}\right)^{2}\right)^{1/4}}\sin(\theta_{\text{SDP}}^{-}-\hat{\delta}(j;m)).$
So for modified Krawtchouk polynomial unbounded double-scaling asymptotic we finally obtain
$\widetilde{K}_{m}(j/n)\bigg{|}_{n=HK,m=gK+s\sqrt{K}+l,K\to\infty}\sim$
$\sim\frac{m!}{n^{m}}\sqrt{\frac{2}{\pi K}}\frac{(\sqrt{1-g})^{m-K-1/2}(\sqrt{1-\mu})^{K/2+j+1/2}}{(\sqrt{g})^{m+1/2}(\sqrt{\mu})^{j-K/2-1/2}}\frac{(1/2)^{m-K/2}\sin(\theta_{\text{SDP}}^{-}-\hat{\delta}(j;m))}{\left((1-g)g-\left(\mu-\frac{1}{2}\right)^{2}\right)^{1/4}},$ (3.11)
where $\theta_{\text{SDP}}^{-}$ and $\hat{\delta}$ are defined above as the functions of $\mu,g$.
3.2. Symplectic polynomials asymptotic
Using the equation (2.9) we can find the asymptotic behavior for $m=gK+s\sqrt{K}+l$,
$n=HK,\;K\sim\infty$ for even $m$
$G_{m}(i/n)=\frac{1}{i^{2}}n^{2}\left(\widetilde{K}_{m+2}(i/n)-\frac{\widetilde{K}_{m+2}(0)}{\widetilde{K}_{m}(0)}\widetilde{K}_{m}(i/n)\right)=$
$=\frac{(m+2)!}{n^{m}i^{2}}\sqrt{\frac{2}{\pi K}}\frac{(\sqrt{1-g})^{m-K+3/2}(\sqrt{1-\mu})^{K/2+i+1/2}}{(\sqrt{g})^{m+5/2}(\sqrt{\mu})^{i-K/2-1/2}}\frac{(1/2)^{m+2-K/2}}{\left((1-g)g-\left(\mu-\frac{1}{2}\right)^{2}\right)^{1/4}}\cdot$
$\cdot\bigg{(}\sin\big{(}\theta_{SDP}^{-}-\hat{\delta}(i;m+2)\big{)}+\frac{(K-m)}{m+2}\frac{g}{1-g}\sin\big{(}\theta_{SDP}^{-}-\hat{\delta}(i;m)\big{)}\bigg{)}+O(1/K).$ (3.12)
4. Limiting behavior of the correlation kernel
The main goal of this section is to compute the asymptotic of the correlation kernel (1.13) in following regime of the symplectic polynomials
$m=gK,\,K=2n+2k,\,n=HK,$ (4.1)
with $K\to\infty$ and fixed $g,H$, and $m$ is the number of polynomial. The equation (3.12) for $m=2n$ has the following form
$G_{2n}(i/n)=g_{2n}(i/n)\sqrt{\Lambda_{2n}}=$
$=\frac{1}{(i/n)^{2}}\frac{(2n+2)!}{n^{2n+2}}F_{2n}(i)\bigg{(}\sin\big{(}\theta_{SDP}^{-}-\hat{\delta}(i;2n+2)\big{)}+\sin\big{(}\theta_{SDP}^{-}-\hat{\delta}(i;2n)\big{)}\bigg{)}+O(1/K),$ (4.2)
where $F_{2n}(i)$ is defined as
$F_{2n}(i)=\frac{(\sqrt{1-2H})^{2n-K+3/2}(\sqrt{1-\mu})^{K/2+i+1/2}}{(\sqrt{2H})^{2n+5/2}(\sqrt{\mu})^{i-K/2-1/2}}\sqrt{\frac{2}{\pi K}}\frac{(1/2)^{2n+2-K/2}}{\left((1-2H)2H-\left(\mu-\frac{1}{2}\right)^{2}\right)^{1/4}},$
with $g=2H$ and in general $20<H<1/$ 2 with $H=1/4$ for $n=k$ (square Young tableaux). Now we substitute asymptotic of the polynomials into the kernel to demonstrate its convergence to the discrete sine-kernel
$\mathcal{K}(u,v)=2\frac{\varkappa_{2n-2}}{\varkappa_{2n}}\frac{g_{2n}(u)g_{2n-2}(v)-g_{2n-2}(u)g_{2n}(v)}{u^{2}-v^{2}}\sqrt{\widetilde{W}(u^{2})\widetilde{W}(v^{2})}=$
$=\frac{(1/2)^{K-1}}{\Lambda_{2n-2}}\frac{G_{2n}(u)G_{2n-2}(v)-G_{2n-2}(u)G_{2n}(v)}{u^{2}-v^{2}}uv\sqrt{\binom{K}{K/2+nu}\binom{K}{K/2+nv}}.$
In fact, substituting $nu=i,vn=j$ and (2.26) we have
$\frac{(1/2)^{K-1}}{\Lambda_{2n-2}}\frac{G_{2n}(u)G_{2n-2}(v)-G_{2n-2}(u)G_{2n}(v)}{u^{2}-v^{2}}uv\sqrt{\binom{K}{K/2+nu}\binom{K}{K/2+nv}}=$
$=\frac{(1/2)^{K-1}\beta_{2n-1}}{S_{2n-1}L_{2n-1}}\frac{F_{2n}(i)F_{2n-2}(j)}{(u^{2}-v^{2})uv}\frac{(2n+2)!(2n!)}{n^{4n+2}}\sqrt{\binom{K}{K/2+i}\binom{K}{K/2+j}}\cdot$
$\cdot\Bigg{(}\bigg{(}\sin\big{(}\theta^{-}_{\text{SDP}}(i;2n+2)-\hat{\delta}(i;2n+2)\big{)}+\sin\big{(}\theta^{-}_{\text{SDP}}(i;2n)-\hat{\delta}(i;2n)\big{)}\bigg{)}\cdot$
$\cdot\bigg{(}\sin\big{(}\theta^{-}_{\text{SDP}}(j;2n)-\hat{\delta}(j;2n)\big{)}+\sin\big{(}\theta^{-}_{\text{SDP}}(j;2n-2)-\hat{\delta}(j;2n-2)\big{)}\bigg{)}-[ij]\bigg{)}\Bigg{)}+O(1/K),$ (4.3)
where $[ij]$ is the last summand with the transposition $i\leftrightarrow j,$ and we use the fact
$\frac{F_{2n-2}(i)F_{2n}(j)}{F_{2n}(i)F_{2n-2}(j)}=1.$
Now using the notation
$\lambda(i;2n)=\theta^{-}_{\text{SDP}}(i;2n)-\hat{\delta}(i;2n).$
we aim to derive the asymptotic of $\lambda$ in the bulk regime for $-1/2<x<1/2$
$u=\frac{1}{H}\left(x+\frac{i^{\prime}}{K}\right),\quad v=\frac{1}{H}\left(x+\frac{j^{\prime}}{K}\right),$
$i=un=xK+i^{\prime}=\frac{K}{2}\bigg{(}1-2\mu(i^{\prime})\bigg{)},\quad j=vn=xK+j^{\prime},$
with $\mu(i^{\prime})=-\frac{1}{2}\left(2x+2i^{\prime}/K-1\right)$. Let us determine the domain of the pieswise functions
$\gamma,\hat{\zeta}_{1},\hat{\zeta}_{2},\hat{\tau},\arg f^{\prime\prime}(t_{0}^{-}))$. We need to choose the interval between the critical points of the action. So we can expand the piecewise functions for example in $g>1/2,\,\mu>1/2$ ($-K/2<i<0$)
$\gamma=\tan^{-1}\left(\frac{\sqrt{(1-g)g-\left(\mu-\frac{1}{2}\right)^{2}}}{1/2-g}\right)+\pi,\quad\zeta_{2}=\tan^{-1}\left(\frac{\sqrt{(1-g)g-\left(\mu-\frac{1}{2}\right)^{2}}}{\mu-\frac{1}{2}}\right),$
$\hat{\zeta}_{1}=\tan^{-1}\left(\frac{\sqrt{g(1-g)-(\mu-1/2)^{2}}}{3/2-g-\mu}\right),\quad\hat{\tau}=\tan^{-1}\left(\frac{\sqrt{(1-g)g-\left(\mu-\frac{1}{2}\right)^{2}}}{g+\mu-1/2}\right),$
$\arg f^{\prime\prime}(t_{0}^{-})=\tan^{-1}\left(\frac{(2g-1)(2\mu-1)}{\sqrt{-4(g-1)g-(1-2\mu)^{2}}}\right)+2\tan^{-1}\left(\frac{-\frac{\mu}{g-1/2}+2g-1}{\sqrt{-4(g-1)g-(1-2\mu)^{2}}}\right)+\pi.$
In the other domains of parameter space such as $g>1/2,\mu<1/2$ we will get the same formulas up to ±$\pm\pi$. These contributions will cancel each other out in final answer. Next we derive the asymptotic of trigonometric part of the kernel accurate up to $O(1/K)$. It is obvious that the shift $m\to m+2$ generate the shift $g\to g+2/K$, so, by the definition of $\lambda(i;2n)$
$\lambda(i;m+2)=\lambda(i;m)-2\zeta_{2}-2\frac{\partial\zeta_{2}}{\partial g}g-\frac{\partial\hat{\zeta}_{1}}{\partial g}2+2\mu\frac{\partial\gamma}{\partial g}+O(1/K).$
Denoting:
$\beta=\left(-2\zeta_{2}-\frac{\partial\hat{\zeta}_{1}}{\partial g}2+\frac{\partial\gamma}{\partial g}\right)\Bigg{|}_{m=2n,i=xK+i^{\prime}}=2\cot^{-1}\left(\frac{x}{\sqrt{2H-4H^{2}-x^{2}}}\right)+O(1/K),$
$\gamma_{i}=\lambda(i;2n),$
we can reduce all trigonometric part of the CD kernel by the simple calculation to:
$\bigg{(}\sin\lambda(i;2n+2)+\sin\lambda(i;2n)\bigg{)}\cdot\bigg{(}\sin\lambda(j;2n)+\sin\lambda(j;2n-2)\bigg{)}-[ij]=$
$=\bigg{(}\sin\gamma_{i}+\sin\gamma_{i}\cos\beta+\cos\gamma_{i}\sin\beta\bigg{)}\bigg{(}\sin\gamma_{j}+\sin\gamma_{j}\cos\beta-\cos\gamma_{j}\sin\beta\bigg{)}-[ij]+O(1/K)=$
$=-2\sin\beta(1+\cos\beta)\sin(\gamma_{i}-\gamma_{j}),$
where:
$-2\sin\beta(1+\cos\beta)=\frac{4x^{3}\sqrt{-g^{2}+g-x^{2}}}{(g-1)^{2}g^{2}}\bigg{(}1+O\left(\frac{1}{K}\right)\bigg{)},$
$\gamma_{i}-\gamma_{j}=-(i^{\prime}-j^{\prime})\left(\pi+\tan^{-1}\left(\frac{2\sqrt{2H-4H^{2}-x^{2}}}{1-4H}\right)\right)\bigg{(}1+O\left(\frac{1}{K}\right)\bigg{)}.$
The normalization of the kernel is not important because it can be cancelled out by the renormalization of the polynomials. Therefore we have established the convergence of the Christoffel–Darboux kernel to the discrete sine-kernel
$\lim\limits_{n\to\infty}\frac{\mathcal{K}\left(\frac{1}{H}\big{(}x+\frac{i^{\prime}}{K}\big{)},\frac{1}{H}\big{(}x+\frac{j^{\prime}}{K}\big{)}\right)}{\mathcal{K}(x/H,x/H)}=\frac{\sin(\pi\rho(x)(i^{\prime}-j^{\prime}))}{\pi\rho(x)(i^{\prime}-j^{\prime})},$ (4.4)
with the limit density
$\rho(x)=\frac{1}{\pi}\cos^{-1}\left(\frac{1-4H}{\sqrt{1-4x^{2}}}\right),$ (4.5)
which ends the proof of Theorem 2. The limit density coincides with the result for the limit shape of symplectic diagrams obtained in [22, Thm. 5.1] after substituting $c=\lim\frac{k}{n}=\frac{1-2H}{2H}$.
In Figure 1 we present the comparison of the correlation kernel from sampling multiple diagrams by Proctor’s algorithm to the discrete sine kernel. We see good agreement of the simulation results with the asymptotic expression even for the relatively small values of $n=50,k=100$.
Figure 1. Plot of the ratio $\frac{\mathcal{K}(i,j)}{\mathcal{K}(i,i)}$ for $i=75$ obtained by sampling $500$ samples, each of $10000$ random Young diagrams inside of $n\times k$ box with $n=50,k=100$ using Proctor’s algorithm (box and whiskers plot) and its comparison to the same ratio for the Christoffel-Darboux kernel (1.9) (dashed black line) and discrete sine kernel (dotted blue) (4.4), (4.5)
5. Conclusion and future work
In this paper, we have investigated the local fluctuations of random Young diagrams arising from the skew Howe duality for symplectic groups $Sp_{2n}\times Sp_{2k}$. The probability measure on diagrams, proportional to the ratio of the dimension of the irreducible $Sp_{2n}\times Sp_{2k}$-representation to the dimension of the exterior algebra $\bigwedge\left(\mathbb{C}^{2n}\otimes\mathbb{C}^{k}\right)$, was expressed as a determinantal point process governed by a correlation kernel derived from semiclassical orthogonal polynomials. Unlike the $GL(n)$ case, where free-fermionic methods apply, the symplectic setting required a Christoffel transformation of Krawtchouk polynomials to construct the relevant orthogonal polynomials. It is an open question how to apply the free-fermionic representation [44] to the skew $Sp_{2n}\times Sp_{2k}$ case. In particular, it is not clear how the tau functions for d-connections can be constructed [45, 46].
By analyzing the integral representation of these polynomials in the double-scaling limit $n,k\to\infty$ with $n/k$ fixed, we have demonstrated that the correlation kernel converges to the discrete sine kernel in the bulk regime. This universality mirrors results for general linear groups [11, 13] but required novel techniques due to the absence of free-fermionic structures.
Our findings align with the limit shape results in [22] and extend the understanding of fluctuations to symplectic symmetry classes. Future work could explore edge asymptotics and connections to discrete Painleve equations. The interplay between representation theory, determinantal processes in classical groups and the discrete integrable systems remains a rich area for further investigation.
6. Appendix
The goal of this appendix is to prove the relation for the second term in the definition (2.9)
$\frac{\widetilde{K}_{m+2}(0)}{\widetilde{K}_{m}(0)}=-\frac{(K-m)(m+1)}{4n^{2}}=\frac{(m+1)(m+2)\binom{K}{m+2}\,_{2}F_{1}\left(-\frac{K}{2},-m-2;-K;2\right)}{4n^{2}\binom{K}{m}\,_{2}F_{1}\left(-\frac{K}{2},-m;-K;2\right)}.$ (6.1)
First of all we need to use the following transformation for $n\in\mathbb{N}$ (see (3.19) in [47])
$\,{}_{2}F_{1}(-n,b,c,z)=\frac{(b)_{n}}{(c)_{n}}(-z)^{n}\,_{2}F_{1}(-n,1-c-n,1-b-n,1/z).$ (6.2)
So we get
$\,{}_{2}F_{1}(-m-2,-K/2,-K,2)=\frac{(-K/2)_{m+2}}{(-K)_{m+2}}(-1)^{m+2}2^{m+2}\,_{2}F_{1}(-m-2,K-m-1,k/2-m-1,1/2).$
Now we can use the standard series for the Gauss hypergeometric function
$\,{}_{2}F_{1}(-m-2,K-m-1,k/2-m-1,1/2)=\sum\limits_{i=0}^{m+2}\frac{(-1)^{i}}{2^{i}}\binom{m+2}{i}\frac{(K-m-1)_{i}}{(K/2-m-1)_{i}}=$
$=\frac{\left((-1)^{m}+1\right)\Gamma\left(\frac{m+3}{2}\right)\Gamma\left(\frac{1}{2}(-K+m+2)\right)}{2\sqrt{\pi}\Gamma\left(-\frac{K}{2}+m+2\right)}=$
$=-\frac{1/4(m+1)(K-m)}{(-K/2+m+1)(-K/2+m)}\,_{2}F_{1}(-m,K-m+1,K/2-m+1,1/2),$
where the last sum can be computed explicitly by the Gauss’s second summation theorem
$\,{}_{2}F_{1}(a,b,1/2(a+b+1),1/2)=\frac{\Gamma(1/2)\Gamma(1/2(a+b+1))}{\Gamma(1/2(1+a))\Gamma(1/2(1+b))}.$
Finally we have
$\,{}_{2}F_{1}(-m-2,-K/2,-K,2)=\frac{(m+1)}{(-K+m+1)}\,_{2}F_{1}(-m,-K/2,-K,2).$
Substituting this into the original expression, we immediately come to the desired proof.